# Principal twistor models and asymptotic hyperkähler metrics

### 04 Mar 2026 v1

@@ 4.3. Future Directions
In this study, we showed that when there exists a hyperkähler metric with asymptotic behavior, the corresponding twistor space can be embedded into the principal twistor model. We consider the following problem on the converse direction:
@@
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and that $X$ admits a crepant resolution $Y$. By Proposition 3.32, the metric $g_{0}$ induces the good triple on $X$. For this good triple, consider the principal twistor model $(\mathcal{Y}(1)\to\mathcal{C}(2),\omega_{P},\tau_{P})$ constructed in Proposition 3.23. In this case, under what conditions is the twistor model $Z_{s}$, obtained by slicing $\mathcal{Y}(1)$ along a real section $s$ of the vector bundle $\mathcal{C}(2)$, actually a twistor space? For example, when the generic fiber of $Z_{s}$ is an affine variety, does $Z_{s}$ admit a family of twistor lines that defines a foliation?
This problem can be reformulated as follows:
@@
The central fiber of the model $Z_{s}$ is diffeomorphic to the crepant resolution $Y$. When $Z_{s}$ admits a family of twistor lines, the following can be said on the asymptotic behavior of the corresponding hyperkähler metric $g_{s}$ on $Y$: Assume that there exists a family of twistor lines on the model $Z_{s}$ that exists on the same sheet as the family of twistor lines corresponding to the cone metric $g_{0}$ with respect to the local covering map $\Phi$ in Proposition 4.5. By tracing the proof of Proposition 4.8 backwards, one can show that the metric $g_{s}$ is asymptotic to the metric $g_{0}$ at infinity on $Y$.
In considering the moduli of hyperkähler structures, we investigated the openness of the twistor space in the principal twistor model (cf. §4.1.3). In this context, the following question remains:
@@
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and that $X$ admits a crepant resolution $Y$. Show that the map $\Psi:\mathcal{M}\to H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$ introduced in Corollary 4.12 is an open map when $X$ has non-isolated singularities.

### 12 Mar 2026 v2

@@ 4.3. Future Directions
In this study, we showed that when there exists a hyperkähler metric with asymptotic behavior, the corresponding twistor space can be embedded into the principal twistor model. We consider the following question in the converse direction:
@@
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and that $X$ admits a crepant resolution $Y$. By Proposition 3.32, the metric $g_{0}$ induces the good triple on $X$. For this good triple, consider the principal twistor model $(\mathcal{Y}(1)\to\mathcal{C}(2),\omega_{P},\tau_{P})$ constructed in Proposition 3.23. In this case, under what conditions is the twistor model $Z_{s}$, obtained by slicing $\mathcal{Y}(1)$ along a real section $s$ of the vector bundle $\mathcal{C}(2)$, actually a twistor space? For example, when the generic fiber of $Z_{s}$ is an affine variety, does $Z_{s}$ admit a family of twistor lines that defines a foliation?
This question can be reformulated as follows:
@@
The central fiber of the model $Z_{s}$ is diffeomorphic to the crepant resolution $Y$. When $Z_{s}$ admits a family of twistor lines, the following can be said on the asymptotic behavior of the corresponding hyperkähler metric $g_{s}$ on $Y$: Assume that there exists a family of twistor lines on the model $Z_{s}$ that exists on the same sheet as those of the cone metric $g_{0}$ with respect to the local covering map $\Phi$ in Proposition 4.5. By tracing the proof of Proposition 4.8 backwards, one can show that the metric $g_{s}$ is asymptotic to the metric $g_{0}$ at infinity on $Y$.
In considering the moduli space of hyperkähler structures, we investigated the openness of the twistor space in the principal twistor model (cf. §4.1.3). In this context, the following question remains:
@@
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and that $X$ admits a crepant resolution $Y$. Show that the map $\Psi:\mathcal{M}\to H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$ introduced in Corollary 4.12 is an open map when $X$ has non-isolated singularities.